OpenStudy (anonymous):

What is the equation of the given line in the point-slope form? MEDAL WILL BE GIVEN

OpenStudy (anonymous):

y-y1 = m(x-x1) (x1,y1) = given points m = slope

OpenStudy (anonymous):

please help @sangya21

OpenStudy (anonymous):

\[slope (m)= \frac{ y_2 - y_1 }{ x_2 - x_1 }\] \[y-y_1 = m(x -x_1)\] (x1y1) = (0,1) (x2,y2) = (3,4) what will be the answer?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[m = \frac{ 4-1 }{ 3 - 0 } = \frac{ 3 }{ 3 } = 1\] now what will be the equation?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

the last what would the equation be?

OpenStudy (anonymous):

what would be your answer?

OpenStudy (anonymous):

y + x = 1?

OpenStudy (anonymous):

@sangya21 ?

OpenStudy (anonymous):

y - y1 = m(x-x1) y-1 = 1(x- 0) y - 1 = x y= x+1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime :)